# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 5.3. The $\widetilde{\Omega}$-action on morphisms
Since the finite Hecke algebra ${\mathcal{H}}_{F}$ is selfinjective, for $F\subseteq\overline{C}$, it follows that every ${\mathcal{H}}_{F}$-module is Gorenstein projective. We may therefore apply 5.20 to $\hat{\omega}$, with $B={\mathcal{H}}$ and $A={\mathcal{H}}_{F}$, to get the result. ∎
Remark 5.22.
Take $A={\mathcal{H}_{\operatorname{aff}}}$ and $B={\mathcal{H}}$ so that $\Omega_{A}=\Omega$, and assume that $\mathbb{G}$ is semisimple and that $p$ does not divide $\left|\Omega\right|$. Then the functor of taking $\Omega$-invariants is exact and the bottom row of (5.9) is exact. Moreover, in that case the two leftmost vertical arrows in (5.9) are isomorphisms and it then follows from the 5-lemma that $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}}\cong[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}_{\operatorname{aff}}}^{\Omega}$ for all ${\mathfrak{m}}_{1},{\mathfrak{m}}_{2}\in{\operatorname{Mod}}({\mathcal{H}})$ with ${\mathfrak{m}}_{1}$ Gorenstein projective. In particular, the functor ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}})$ is faithful in that case. In general though, letting $\mathfrak{c}=\ker(\theta)$ and recalling that $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}}\cong{\operatorname{Ext}}^{1}_{\mathcal{H}}({\mathfrak{m}}_{1},\mathfrak{c})$ and $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{\mathcal{H}_{\operatorname{aff}}}^{\Omega}\cong{\operatorname{Ext}}^{1}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}_{1},\mathfrak{c})^{\Omega}$ (c.f. 2.16), the exact sequence in [3, §4.2, eq.(4.4)] shows that the kernel of the map $[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{B}\to[{\mathfrak{m}}_{1},{\mathfrak{m}}_{2}]_{A}^{\Omega_{A}}$ is $H^{1}(\Omega,{\operatorname{Hom}}_{\mathcal{H}_{\operatorname{aff}}}({\mathfrak{m}}_{1},\mathfrak{c}))$ and hence may fail to be zero when $\Omega$ has non-trivial group cohomology. In general, ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{\mathcal{H}_{\operatorname{aff}}})$ may thus fail to be faithful.
We now apply the above construction to study isomorphism classes of simple supersingular modules in ${\operatorname{Ho}}({\mathcal{H}})$. We let $\chi$ be a supersingular character of ${\mathcal{H}_{\operatorname{aff}}}$ and consider two irreducible, finite dimensional representations $V$ and $V^{\prime}$ of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$. Let ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}={\mathcal{H}}\otimes_{{\mathcal{H}}_{\chi}}(\chi\otimes V^{\prime})$, and fix $g\in[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}}$.
First, observe that for any $\mathfrak{n}\in{\operatorname{Mod}}({\mathcal{H}}_{\chi})$ we obtain from the braid relations a direct sum decomposition
$\mathfrak{n}\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}\cong\bigoplus_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}\mathfrak{n}\otimes T_{\tilde{\omega}}.$
This is in fact a decomposition as an ${\mathcal{H}}_{\chi}$-module. Indeed, for all $h\in{\mathcal{H}_{\operatorname{aff}}}$ and all $x\in\mathfrak{n}$ we have $(x\otimes T_{\tilde{\omega}})\cdot h=x(T_{\tilde{\omega}}hT_{\tilde{\omega}}^{-1})\otimes T_{\tilde{\omega}}$. Furthermore, the group $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{\chi^{\tilde{\omega}}}=\tilde{\omega}^{-1}\widetilde{\Omega}_{\chi}\tilde{\omega}$ naturally acts on $T_{\tilde{\omega}}\otimes\mathfrak{n}$. Note that these actions do not depend, up to isomorphism, on the choice of coset representatives. When $\mathfrak{n}=\chi\otimes V^{\prime}$, the summand $\mathfrak{n}\otimes T_{\tilde{\omega}}$ naturally identifies with $\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}$ as an ${\mathcal{H}}_{\chi}$-module. We therefore get a decomposition
(5.11) $({\mathfrak{m}}^{\prime})|_{{\mathcal{H}}_{\chi}}\cong\bigoplus_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}},$
and similarly for ${\mathfrak{m}}|_{{\mathcal{H}}_{\chi}}$. Correspondingly, we have ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{\chi}})(g)=(g_{\tilde{\omega}})_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}$, where $g_{\tilde{\omega}}\in[{\mathfrak{m}},\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}]_{{\mathcal{H}}_{\chi}}$.
@@
For $\tilde{\omega}\in\widetilde{\Omega}/\widetilde{\Omega}_{\chi}$, we let $\iota_{\tilde{\omega}}:\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}}\to{\mathfrak{m}}$ denote the natural ${\mathcal{H}}_{\chi}$-linear inclusion. Furthermore, let $l$ denote the index of $\widetilde{\Omega}_{\chi}$ in $\widetilde{\Omega}$. From (5.11), we may then regard elements of $[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{{\mathcal{H}}_{F}}\cong\bigoplus_{\tilde{\omega},\tilde{\omega}^{\prime}}[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}^{\prime}}\otimes(V^{\prime})^{\tilde{\omega}^{\prime}}]_{{\mathcal{H}}_{F}}$ as $l\times l$ matrices with entries in the Hom spaces $[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}^{\prime}}\otimes(V^{\prime})^{\tilde{\omega}^{\prime}}]_{{\mathcal{H}}_{F}}$, and the image of any $g^{\prime}\in[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}}$ under ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})$ is the matrix $({\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})(g^{\prime}_{\tilde{\omega}^{\prime}}\circ\iota_{\tilde{\omega}}))_{\tilde{\omega},\tilde{\omega}^{\prime}}$. Since the composite $g^{\prime}_{\tilde{\omega}^{\prime}}\circ(\iota_{\tilde{\omega}})$ has image under ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})$ equal to zero for $\tilde{\omega}\neq\tilde{\omega}^{\prime}$ by 5.7(ii) and 5.8, we deduce that ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(g^{\prime})$ is in fact a diagonal matrix.
Applying the above to $g^{\prime}=g$, we deduce that the diagonal matrix ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(g)$ is an isomorphism and thus for all $\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}$ we have that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})(g_{\tilde{\omega}}\circ\iota_{\tilde{\omega}})$ is an isomorphism in ${\operatorname{Ho}}({\mathcal{H}}_{F})$ as well. By applying 5.20 with $A={\mathcal{H}}_{F}$, $B={\mathcal{H}}_{\chi}$ and noting that $\widetilde{\Omega}_{A}=\widetilde{\Omega}_{F,\chi}$ in that case, we see that this isomorphism is in fact in $[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}]^{\widetilde{\Omega}_{F,\chi}}_{{\mathcal{H}}_{F}}$. Finally, by applying 5.6 we see that this isomorphism is an actual $\widetilde{\Omega}_{F,\chi}$-equivariant isomorphism in the category ${\operatorname{Mod}}({\mathcal{H}}_{F})$ whenever $(\chi^{\tilde{\omega}})|_{{\mathcal{H}}_{F}}$ is not projective, as required. ∎
@@
Our assumptions say precisely that we are not in the setting of 5.7(i). We may thus apply 5.17 to deduce that $\chi$ and $\chi^{\prime}$ are $\widetilde{\Omega}$-conjugate. By conjugating the pair $(\chi^{\prime},V^{\prime})$, we may therefore assume that $\chi=\chi^{\prime}$. We can then apply 5.23 to the face $F\subseteq\overline{C}$ given in the statement to deduce that $V\cong V^{\prime}$ as representations of $\widetilde{\Omega}_{F,\chi}=\widetilde{\Omega}_{\chi}$. This shows that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$. ∎
In general, it may be that $\widetilde{\Omega}_{F,\chi}\neq\widetilde{\Omega}_{\chi}$ whenever $\chi|_{{\mathcal{H}}_{F}}$ is not projective so that the above Theorem does not apply (see e.g. 5.31). However, our techniques still apply in many situations. To illustrate this, we finish this paper by classifying isomorphism classes of simple supersingular ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$ for $G$ an arbitrary finite product of general linear groups.
@@ 5.4. Products of $\mathrm{GL}_{n}$’s
We make a few preliminary remarks. First note that $\hat{\omega}_{i}^{n_{i}}={\operatorname{diag}}(\varpi,\ldots,\varpi)$ is central in $\widetilde{W}$. It follows that $\hat{\omega}_{1}^{n_{1}},\ldots,\hat{\omega}_{r}^{n_{r}}$ and $\widetilde{\Omega}_{T^{\prime}}$ generate together an index $n_{1}\cdots n_{r}$ subgroup of $\widetilde{\Omega}$ which stabilises every character of ${\mathcal{H}_{\operatorname{aff}}}$. If we fix such a character $\chi$, it follows that we have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{d_{1}}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{d_{r}}\rangle)$ for some $d_{1},\ldots,d_{r}\geq 1$ such that $d_{i}$ divides $n_{i}$ for all $1\leq i\leq r$. If $V$ is any finite dimensional representation of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$, we then have that $\hat{\omega}_{1}^{d_{1}},\ldots,\hat{\omega}_{r}^{d_{r}}$ and the generators of $\Omega_{T^{\prime}}$ have a simultaneous eigenvector (since they commute) and this eigenvector spans a one dimensional subrepresentation. If $V$ is irreducible it must therefore be one dimensional. When $\chi$ is supersingular the corresponding simple supersingular ${\mathcal{H}}$-module $(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ then has $k$-dimension $d_{1}\cdots d_{r}$.
@@
    If there exists $s,s^{\prime}\in S_{i}$ adjacent in the affine Dynkin diagram, for some $1\leq i\leq r$, such that $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$ then $d_{i}>1$.
@@
    If $S\neq S_{\xi}$ then $d_{i}>1$ for every $i$ such that $S_{i}\setminus S_{\xi}\neq\emptyset$.
@@
For (i), since the action of $\omega_{i}$ permutes the nodes of the affine Dynkin diagram of $\mathrm{GL}_{n_{i}}$ cyclically, c.f. [16, §1.8], without loss of generality we may assume $s^{\prime}=\omega_{i}s\omega_{i}^{-1}$. Thus we have that $\chi(T_{\hat{s}})\neq\chi^{\hat{\omega}_{i}}(T_{\hat{s}})$ and hence $\hat{\omega}_{i}\notin\widetilde{\Omega}_{\chi}$ as required.
For (ii), suppose that we have $s\in S_{i}\setminus S_{\xi}$. Then there is some $1\leq j\leq n_{i}$ such that the subtorus $T_{s}(\mathbb{F}_{q})$ of $T_{i}(\mathbb{F}_{q})$ is equal to $\{{\operatorname{diag}}(t)_{j}\cdot{\operatorname{diag}}(t^{-1})_{j+1}\mid t\in\mathbb{F}_{q}\}$, where the subscripts are taken modulo $n_{i}$. The condition that $s\notin S_{\xi}$ then corresponds to saying that there exists $t\in\mathbb{F}_{q}$ such that $\xi({\operatorname{diag}}(t)_{j})\neq\xi({\operatorname{diag}}(t)_{j+1})$. Since ${\operatorname{diag}}(t)_{j+1}=\hat{\omega}_{i}{\operatorname{diag}}(t)_{j}\hat{\omega}_{i}^{-1}$, this shows that $\xi^{\hat{\omega}_{i}}\neq\xi$ and thus $\hat{\omega}_{i}\notin\widetilde{\Omega}_{\chi}$ as required.
@@
We write $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{d_{1}}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{d_{r}}\rangle)$ and we pick $s\in S\setminus S_{\xi}$, say with $s\in S_{i}$. Note that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective for any $F\subseteq\overline{C}$ such that $s\in S_{F}$ by 5.10(i). In particular, this applies to the face $F$ given by $S_{F}=\{\omega_{i}^{-d_{i}l}s\omega_{i}^{d_{i}l}\mid 0\leq l<n_{i}/d_{i}\}$, which is well-defined by 5.25(ii). Also note that this $F$ satisfies $\Omega_{F}=\langle\omega_{i}^{d_{i}}\rangle\times\Omega_{T^{\prime}}\times\prod_{j\neq i}\Omega_{j}$ and thus we have $\widetilde{\Omega}_{\chi}\subseteq\widetilde{\Omega}_{F}$. We may therefore apply 5.24 to obtain the result. ∎
@@
When $\mathbb{G}$ is a reductive group whose root system is irreducible and not of type $A$, then the action of $\Omega$ on the affine Dynkin diagram can never be transitive except in type $D_{4}$ (c.f. [16, §1.8]). Hence, in those cases, we also have that two nonisomorphic simple supersingular ${\mathcal{H}}$-modules are still nonisomorphic in ${\operatorname{Ho}}({\mathcal{H}})$ whenever they containing a common supersingular character $\chi=(J,\xi)$ of ${\mathcal{H}_{\operatorname{aff}}}$ such that $S_{\xi}\neq S$. The argument is the same as the above proof: if we pick $s\in S\setminus S_{\xi}$ then the face $F$ such that $S_{F}$ is equal to the orbit of $s$ under $\Omega_{\chi}$, which is well-defined by the non-transitivity of the $\Omega$-action, allows us to conclude using 5.24.
In the case where $S=S_{\xi}$, the following result will be key.

### v4

@@ 5.3. The $\widetilde{\Omega}$-action on morphisms
Apply 5.22 to $\hat{\omega}$, with $B={\mathcal{H}}$ and $A={\mathcal{H}}_{F}$, to get the result. ∎
We now apply the above construction to study isomorphism classes of simple supersingular modules in ${\operatorname{Ho}}({\mathcal{H}})$. We let $\chi$ be a supersingular character of ${\mathcal{H}_{\operatorname{aff}}}$ and consider two irreducible, finite dimensional representations $V$ and $V^{\prime}$ of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$. Let ${\mathfrak{m}}=(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ and ${\mathfrak{m}}^{\prime}=(\chi\otimes V^{\prime})\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$, and fix $g\in[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}}$. Recall that we have a decomposition $({\mathfrak{m}}^{\prime})|_{{\mathcal{H}}_{\chi}}\cong\bigoplus_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}$, and similarly for ${\mathfrak{m}}|_{{\mathcal{H}}_{\chi}}$ (cf. (5.4)). Correspondingly, we have ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{\chi}})(g)=(g_{\tilde{\omega}})_{\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}}$, where $g_{\tilde{\omega}}\in[{\mathfrak{m}},\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}]_{{\mathcal{H}}_{\chi}}$.
@@
For $\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}$, we let $\iota_{\tilde{\omega}}:\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}}\to{\mathfrak{m}}$ denote the natural ${\mathcal{H}}_{\chi}$-linear inclusion. Furthermore, let $l$ denote the index of $\widetilde{\Omega}_{\chi}$ in $\widetilde{\Omega}$. By (5.4), we may then regard elements of $[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{{\mathcal{H}}_{F}}\cong\bigoplus_{\tilde{\omega},\tilde{\omega}^{\prime}}[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}^{\prime}}\otimes(V^{\prime})^{\tilde{\omega}^{\prime}}]_{{\mathcal{H}}_{F}}$ as $l\times l$ matrices with entries in the Hom spaces $[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}^{\prime}}\otimes(V^{\prime})^{\tilde{\omega}^{\prime}}]_{{\mathcal{H}}_{F}}$, and the image of any $g^{\prime}\in[{\mathfrak{m}},{\mathfrak{m}}^{\prime}]_{\mathcal{H}}$ under ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})$ is the matrix $({\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})(g^{\prime}_{\tilde{\omega}^{\prime}}\circ\iota_{\tilde{\omega}}))_{\tilde{\omega},\tilde{\omega}^{\prime}}$. Since the composite $g^{\prime}_{\tilde{\omega}^{\prime}}\circ\iota_{\tilde{\omega}}$ has image under ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})$ equal to zero for $\tilde{\omega}\neq\tilde{\omega}^{\prime}$ by 5.8(ii) and 5.9, we deduce that ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(g^{\prime})$ is in fact a diagonal matrix.
Applying the above to $g^{\prime}=g$, we deduce that the diagonal matrix ${\operatorname{Ho}}({\operatorname{Res}}^{\mathcal{H}}_{{\mathcal{H}}_{F}})(g)$ is an isomorphism and thus for all $\tilde{\omega}\in\widetilde{\Omega}_{\chi}\backslash\widetilde{\Omega}$ we have that ${\operatorname{Ho}}({\operatorname{Res}}^{{\mathcal{H}}_{\chi}}_{{\mathcal{H}}_{F}})(g_{\tilde{\omega}}\circ\iota_{\tilde{\omega}})$ is an isomorphism in ${\operatorname{Ho}}({\mathcal{H}}_{F})$ as well. By applying 5.22 with $A={\mathcal{H}}_{F}$, $B={\mathcal{H}}_{\chi}$ and noting that $\widetilde{\Omega}_{A}=\widetilde{\Omega}_{F,\chi}$ in that case, we see that this isomorphism is in fact in $[\chi^{\tilde{\omega}}\otimes V^{\tilde{\omega}},\chi^{\tilde{\omega}}\otimes(V^{\prime})^{\tilde{\omega}}]^{\widetilde{\Omega}_{F,\chi}}_{{\mathcal{H}}_{F}}$. Finally, by applying 5.7 we see that this isomorphism is an actual $\widetilde{\Omega}_{F,\chi}$-equivariant isomorphism in the category ${\operatorname{Mod}}({\mathcal{H}}_{F})$ whenever $(\chi^{\tilde{\omega}})|_{{\mathcal{H}}_{F}}$ is not projective, as required. ∎
@@
Our assumptions say precisely that we are not in the setting of 5.8(i). We may thus apply 5.19 to deduce that $\chi$ and $\chi^{\prime}$ are $\widetilde{\Omega}$-conjugate. By conjugating the pair $(\chi^{\prime},V^{\prime})$, we may therefore assume that $\chi=\chi^{\prime}$. We can then apply 5.25 to the face $F\subseteq\overline{C}$ given in the statement to deduce that $V\cong V^{\prime}$ as representations of $\widetilde{\Omega}_{F,\chi}=\widetilde{\Omega}_{\chi}$. This shows that ${\mathfrak{m}}\cong{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}({\mathcal{H}})$. ∎
In general, it may be that $\widetilde{\Omega}_{F,\chi}\neq\widetilde{\Omega}_{\chi}$ whenever $\chi|_{{\mathcal{H}}_{F}}$ is not projective so that the above Theorem does not apply (see e.g. 5.34). However, our techniques still apply in many situations. To illustrate this, we finish this paper by classifying isomorphism classes of simple supersingular ${\mathcal{H}}$-modules in ${\operatorname{Ho}}({\mathcal{H}})$ for $G$ an arbitrary finite product of general linear groups.
@@ 5.4. Products of $\mathrm{GL}_{n}$’s
We make a few preliminary remarks. In order to study simple supersingular ${\mathcal{H}}$-modules, we need to consider the stabiliser $\widetilde{\Omega}_{\chi}$ of a supersingular character.
Lemma 5.27.
Let $\chi$ be a character of ${\mathcal{H}_{\operatorname{aff}}}$. Then we have $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{d_{1}}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{d_{r}}\rangle)$ for some $d_{1},\ldots,d_{r}\geq 1$ such that $d_{i}$ divides $n_{i}$ for all $1\leq i\leq r$.
Proof.
For each $1\leq i\leq r$, we write ${\mathcal{H}}_{{\operatorname{aff}},i}$ for the $k$-subalgebra of ${\mathcal{H}_{\operatorname{aff}}}$ generated by $\{T_{\hat{s}}\mid s\in S_{i}\}$ and $T_{i}(\mathbb{F}_{q})$. Note that ${\mathcal{H}_{\operatorname{aff}}}\cong{\mathcal{H}}_{{\operatorname{aff}},1}\otimes_{k}\ldots\otimes_{k}{\mathcal{H}}_{{\operatorname{aff}},r}$ as $k$-algebras, and that $\widetilde{\Omega}$ normalises each ${\mathcal{H}}_{{\operatorname{aff}},i}$. Correspondingly, $\chi\cong\chi_{1}\otimes_{k}\cdots\otimes_{k}\chi_{r}$ where $\chi_{i}:=\chi|_{{\mathcal{H}}_{{\operatorname{aff}},i}}$ ($1\leq i\leq r$). We fix some $1\leq i\leq r$. The action of $\widetilde{\Omega}$ on characters of ${\mathcal{H}}_{{\operatorname{aff}},i}$ factors through $\Omega_{i}=\langle\omega_{i}\rangle\cong{\mathbb{Z}}$. Since $\hat{\omega}_{i}^{n_{i}}={\operatorname{diag}}(\varpi,\ldots,\varpi)$ is central in $\widetilde{W}$, we deduce that the stabiliser $\widetilde{\Omega}_{\chi_{i}}$ equals $\widetilde{\Omega}_{T^{\prime}}\rtimes\langle\hat{\omega}_{i}^{d_{i}}\rangle\times\prod_{j\neq i}\langle\hat{\omega}_{j}\rangle$ for some $d_{i}|n_{i}$. Since $\widetilde{\Omega}_{\chi}=\bigcap_{1\leq i\leq r}\widetilde{\Omega}_{\chi_{i}}$, we deduce the claimed result. ∎
We continue to consider a character $\chi$ of ${\mathcal{H}_{\operatorname{aff}}}$. If $V$ is any finite dimensional representation of $\widetilde{\Omega}_{\chi}$ on which $T(\mathbb{F}_{q})$ acts by $\chi$, we then have that $\hat{\omega}_{1}^{d_{1}},\ldots,\hat{\omega}_{r}^{d_{r}}$ (with $d_{i}$ as in 5.27) and the generators of $\Omega_{T^{\prime}}$ have a simultaneous eigenvector (since they commute) and this eigenvector spans a one dimensional subrepresentation. If $V$ is irreducible it must therefore be one dimensional. When $\chi$ is supersingular the corresponding simple supersingular ${\mathcal{H}}$-module $(\chi\otimes V)\otimes_{{\mathcal{H}}_{\chi}}{\mathcal{H}}$ then has $k$-dimension $d_{1}\cdots d_{r}$.
@@
    If there exist $s,s^{\prime}\in S_{i}$ adjacent in the affine Dynkin diagram, for some $1\leq i\leq r$, such that $\chi(T_{\hat{s}})\neq\chi(T_{\hat{s^{\prime}}})$ then $d_{i}>1$.
@@
    If $S\neq S_{\xi}$ then $d_{i}>1$ for every $i$ such that $S_{i}\nsubseteq S_{\xi}$.
@@
For (i), since the action of $\omega_{i}$ permutes the nodes of the affine Dynkin diagram of $\mathrm{GL}_{n_{i}}$ cyclically, cf. [18, §1.8], without loss of generality we may assume $s^{\prime}=\omega_{i}s\omega_{i}^{-1}$. Thus we have that $\chi(T_{\hat{s}})\neq\chi^{\hat{\omega}_{i}}(T_{\hat{s}})$ and hence $\hat{\omega}_{i}\notin\widetilde{\Omega}_{\chi}$ as required.
For (ii), suppose that we have $s=s_{\alpha}\in S_{i}\setminus S_{\xi}$. Then there is some $1\leq j\leq n_{i}$ such that the subtorus $\alpha^{\vee}(\mathbb{F}_{q}^{\times})$ of $T_{i}(\mathbb{F}_{q})$ is equal to $\{{\operatorname{diag}}(t)_{j}\cdot{\operatorname{diag}}(t^{-1})_{j+1}\mid t\in\mathbb{F}_{q}\}$, where the subscripts are taken modulo $n_{i}$. The condition that $s\notin S_{\xi}$ then corresponds to saying that there exists $t\in\mathbb{F}_{q}$ such that $\xi({\operatorname{diag}}(t)_{j})\neq\xi({\operatorname{diag}}(t)_{j+1})$. Since ${\operatorname{diag}}(t)_{j+1}=\hat{\omega}_{i}{\operatorname{diag}}(t)_{j}\hat{\omega}_{i}^{-1}$, this shows that $\xi^{\hat{\omega}_{i}}\neq\xi$ and thus $\hat{\omega}_{i}\notin\widetilde{\Omega}_{\chi}$ as required.
@@
We write $\widetilde{\Omega}_{\chi}=\widetilde{\Omega}_{T^{\prime}}\rtimes(\langle\hat{\omega}_{1}^{d_{1}}\rangle\times\cdots\times\langle\hat{\omega}_{r}^{d_{r}}\rangle)$ and we pick $s\in S\setminus S_{\xi}$, say with $s\in S_{i}$. Note that ${\operatorname{Res}}^{\mathcal{H}_{\operatorname{aff}}}_{{\mathcal{H}}_{F}}(\chi)$ is not projective for any $F\subseteq\overline{C}$ such that $s\in S_{F}$ by 5.12(i). In particular, this applies to the face $F$ given by $S_{F}=\{\omega_{i}^{-d_{i}l}s\omega_{i}^{d_{i}l}\mid 0\leq l<n_{i}/d_{i}\}$, which is well-defined by 5.28(ii). Also note that this $F$ satisfies $\Omega_{F}=\langle\omega_{i}^{d_{i}}\rangle\times\Omega_{T^{\prime}}\times\prod_{j\neq i}\Omega_{j}$ and thus we have $\widetilde{\Omega}_{\chi}\subseteq\widetilde{\Omega}_{F}$. We may therefore apply 5.26 to obtain the result. ∎
@@
When $\mathbb{G}$ is a reductive group whose root system is irreducible and not of type $A$, then the action of $\Omega$ on the affine Dynkin diagram can never be transitive (cf. [18, §1.8]). Hence, in those cases, we also have that two nonisomorphic simple supersingular ${\mathcal{H}}$-modules are still nonisomorphic in ${\operatorname{Ho}}({\mathcal{H}})$ whenever they contain a common supersingular character $\chi=(J,\xi)$ of ${\mathcal{H}_{\operatorname{aff}}}$ such that $S_{\xi}\neq S$. The argument is the same as the above proof: if we pick $s\in S\setminus S_{\xi}$ then the face $F$ such that $S_{F}$ is equal to the orbit of $s$ under $\Omega_{\chi}$, which is well-defined by the non-transitivity of the $\Omega$-action, allows us to conclude using 5.26.
In the case where $S=S_{\xi}$, the following result will be key in order to apply 5.26.